# Variance-Reducing Couplings for Random Features: Perspectives from Optimal Transport

### v1

@@ Appendix D Asymptotic optimality of permutation densities for variance reduction
In Sec. 4.1 of the main text, we alluded to the choice of permutation densities in Eq. 14 being not only convenient (in terms of ease of sampling and optimising), but also well-motivated by results from OT theory. Here, we make this statement concrete, showing that in the limit of infinite permutation order $n$ this class contains the solution to a broad range of OT problems.
@@
Consider the class of joint measures $\Lambda^{(\sigma)}\coloneqq\{\mu^{(\sigma)}\}$ specified by permutations $\sigma$ of order $n$, given by permutation densities from Eq. 14 pushed forward coordinate-wise using the (left-continuous) inverse CDF $F_{\eta}^{-1}$. Suppose also that the marginal measure $\eta$ is absolutely continuous with respect to Lebesgue measure. Consider also a continuous function $f:\mathbb{R}\to\mathbb{R}$ whose expectation is to be estimated using $m=2$ coupled samples. In the limit $n\to\infty$, the class $\Lambda^{(\sigma)}$ contains measures that converge weakly to the optimal transport plan – that is, the sample coupling which provides the smallest possible estimator variance.
@@
$\mu^{*}=\textrm{arg inf}_{\mu\in\Lambda_{2}(\eta)}\left[\mathbb{E}_{(\omega_{1},\omega_{2})\sim\mu}f(\omega_{1})f(\omega_{2})\right].$ (61)
@@
$\mu_{n}^{*}=\textrm{arg inf}_{\mu\in\Lambda_{2}(\eta_{n})}\left[\mathbb{E}_{(\omega_{1},\omega_{2})\sim\mu}f(\omega_{1})f(\omega_{2})\right].$ (62)
@@
The OT plan is (asymptotically) in our search class. Lastly, we have that our search class of couplings $\Lambda^{(\sigma)}$ (corresponding to permutation densities from Eq. 14 pushed forward by $F_{\eta}^{-1}$) contains measures that converge in distribution to $\mu_{n}^{(\sigma)}$ when $n\to\infty$. In this mathematical sense, the asymptotic limit the class of couplings amongst which we optimise includes measures that give the greatest possible variance reduction: more glibly, our method is ‘asymptotically optimal’. This is intuitive because as the order of the permutation grows each ‘tile’ of nonzero density narrows and can be increasingly well-approximated by a delta function.
@@ E.1 Setup: estimating PageRank with random walks
$\widetilde{\mathbf{P}}\coloneqq(1-p)\mathbf{P}+\frac{p}{N}\mathbf{E}.$ (67)
Here, $p\in(0,1)$ is a scalar, $N$ is the number of nodes and $\mathbf{E}=[1]_{i,j\in\mathcal{N}}$ is a matrix whose entries are all ones. $\mathbf{P}$ is the transiton matrix of a simple random walk,
@@
where we normalised the sum of vector entries to $1$. Computing $\bm{\rho}$ analytically is expensive. Rearranging and Taylor expanding $(1-(1-p)\mathbf{P})^{-1}$, it is straightforward to see that
$\bm{\rho}_{i}=\frac{p}{N}\sum_{j\in\mathcal{N}}\sum_{k=0}^{\infty}(1-p)^{k}\mathbf{P}^{k}_{ji}.$ (70)
This is simply a sum over all walks from each of the graph nodes $v_{j}$ to node $v_{i}$, weighted by their respective probabilities – that is, the expected number of random walkers ending at node $v_{i}$ if they terminate with probability $p$ at every timestep – which invites the estimator proposed by Fogaras et al. [2005],
@@ E.2 OT formulation of PageRank variance reduction
[…]
@@ E.3 Empirical results for PageRank
Figure 8: PageRank estimator errors $\|\bm{\rho}-\widehat{\bm{\rho}}\|_{2}$ for a range of real-world graphs when $2$ random walkers are taken to be i) i.i.d., ii) antithetic or iii) $\sigma$-coupled. Lower is better. The error is always at least as small and often substantially better when the walkers are coupled using our OT coupling. Estimates of one standard deviation are shaded.
We now provide experiments using our OT coupling to improve the convergence of PageRank estimates. Fig. 8 shows the results. For termination probabilities $p\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$, we compute the optimal permutation $\sigma$ by solving the matching problem in Eq. 74 on an Erdős-Rènyi graph with $N=100$ nodes, taking permutation order $n=10$ as a hyperparameter. We then test these couplings on a range of real-world graphs [Ivashkin, 2023] of different sizes, plotting the estimator error $\|\bm{\rho}-\widehat{\bm{\rho}}\|_{2}$ (where $\bm{\rho}$ is the true PageRank vector and $\widehat{\bm{\rho}}$ is its MC estimate) against the termination probability $p$. We include walkers that are i.i.d., antithetic [Reid et al., 2024c] and $\sigma$-coupled. At every value of $p$, our method performs at least as well as the baselines, and in some cases (e.g. on the very large cora graph) it gives much greater variance reduction.
Note that we only consider connected subgraphs and take $1000$ repeats to estimate the standard deviations. Reading the plots from left to right for the top row and then the bottom, the number of nodes are: $\{34,62,105,115,398,986,1133,1272,2120,2708\}$. The shape of the curve and size of the gain from coupling depends on the particular graph being considered. For some graphs it is not possible to obtain a big reduction in PageRank estimator variance by coupling walkers (e.g. football and newsgroup), so neither antithetic termination nor our $\sigma$-coupling can provide a big gain. A detailed investigation of how these observations relate to the mathematical graph structure is deferred to future work.
@@ F.1 More results for Gram matrix estimation
First, we run the experiment in the first part of Sec. 4.2 for more real-world graphs. We approximate the $2$-regularised Laplacian kernel $\smash{(\mathbf{I}-\sigma^{2}\widetilde{\mathbf{L}})^{-2}}$ with a regularisation parameter $\sigma=1$ and $m=2$ walkers. Fig. 9 shows the results, with the kernel approximation error normalised by the i.i.d. result (unlike in Fig. 1) for clarity of comparison. The $\sigma$-coupling consistently reduces the estimator variance, providing equally good or better results compared to the hand-crafted antithetic termination algorithm. The size of the gain depends on the graph structure. More work is needed to characterise the type of graph for which coupling walk lengths gives the biggest reductions in kernel estimator error.
Figure 9: Relative Frobenius norm error $\|\mathbf{K}-\widehat{\mathbf{K}}\|_{\textrm{F}}/\|\mathbf{K}\|_{\textrm{F}}$ between true and approximated Gram matrices for different termination probabilities when walkers are i) i.i.d., ii) antithetic [Reid et al., 2024c] or iii) $\sigma$-coupled using our OT approach. All results are normalised by the i.i.d. variance for easy comparison. Lower is better. The $\sigma$-coupling is consistently as good or better across values of $p$ and graph topologies and sizes. Standard deviation estimates are shaded.
@@ F.2 Scalable GPs on graphs
Like their Euclidean counterparts, exact GPs on graphs suffer $\mathcal{O}(N^{3})$ time complexity, making them impractical for very large structures. Techniques to improve their scalability include graph Fourier feature approximations, which approximate the kernel matrix using a truncated eigenvalue expansion computed using the Lanczos algorithm.  The price of increased speed is that our kernel estimate is biased. Another approach is to restrict to kernels whose inverses are sparse, e.g. a Matérn kernel with a small integer $\nu$-parameter, at the expense of lost flexibility [Borovitskiy et al., 2021]. In this paper, we have proposed to instead use GRFs, which give an unbiased estimate to the kernel matrix in subquadratic time. GRFs are sparse: intuitively, $\phi_{\textrm{GRF}}(v_{i})$ is constructed by simulating $m$ random walks out of $v_{i}$ and adding weighted contributions only at the coordinates corresponding to visited nodes. For walks with geometrically-distributed lengths this is typically a small subset of $\mathcal{N}$. Therefore, the matrix $\smash{\widehat{\mathbf{K}}\coloneqq[\phi(v_{i})^{\top}\phi(v_{j})]_{i,j=1}^{N}}$ is a sparse, unbiased estimate to any kernel, not just specific families. [^5] Strictly: any graph node kernel that is expressible as a function of a weighted adjacency matrix. We can use established numerical routines for sparse linear algebra to speed up training, providing computational savings on big graphs. This is analagous to the sparse spectrum GPs discussed in App. B [Lázaro-Gredilla et al., 2010]; the interested reader might benefit from reviewing this section first.
@@
Training GRF-GPs: technical details. Here, provide details to supplement Sec. 4.2. We use mesh graphs made available by Dawson-Haggerty [2023], taking the faces as our graph nodes. Reading left to right, the number of nodes are: $894,999,1280,1572,3838,8700$. For probabilistic mesh interpolation, we use the diffusion kernel $\mathbf{K}=\kappa\exp(-\gamma^{2}\widetilde{\mathbf{L}})$. Here, $\widetilde{\mathbf{L}}$ is the normalised graph Laplacian (Eq. 40) and $\kappa,\gamma$ are learnable parameters corresponding to the signal variance and kernel lengthscale respectively. The observation noise $\sigma$ is also learnable. For every graph, we construct a kernel estimate by sampling $m=8$ i.i.d. random walks out of every node. We then train the corresponding vanilla GP by optimising the log marginal likelihood on observed nodes, using the Adam optimiser [Kingma and Ba, 2014] for $1000$ epochs (after which both the objective value and the hyperparameters are empirically found to have converged). We freeze $\kappa,\gamma,\sigma$ and compute the corresponding exact kernel. We approximate this exact kernel using $\{16,32,64\}$ random walks that are either i.i.d., antithetic or $\sigma$-coupled. We perform approximate Bayesian inference in each case, computing the accuracy of the kernel estimate, the average test RMSE and the KL divergence to the true posterior distribution (unnormalised by the number of test points). We use permutations of order $n=50$ and a termination probability $p=0.4$. As reported in the main body, coupling the lengths of graph random walkers permits better approximate Bayesian inference with GPs on graphs.
@@ F.3 Probabilistic traffic interpolation
Figure 10: Results for probabilistic traffic interpolation experiment. Plots show the kernel approximation accuracy, test RMSE and KL divergence to the true posterior for graph kernels estimated using i.i.d., antithetic and $\sigma$-coupled GRFs. Lower is better. $\sigma$-coupling gives the best results for approximate inference. Estimates of the standard deviation over random draws of walkers are shaded.

### v2

@@ Appendix D Asymptotic optimality of permutation densities for variance reduction
In Sec. 4.1 of the main text, we alluded to the choice of permutation densities $p_{\sigma}(x,y)\coloneqq n1_{\sigma(\lceil nx\rceil)=\lceil ny\rceil}$ depicted in Fig. 2 being not only convenient (in terms of ease of sampling and optimising), but also well-motivated by results from OT theory. Here, we make this statement concrete, showing that in the limit of infinite permutation order $n$ this class contains the solution to a broad range of OT problems.
@@
Consider the class of joint measures $\Lambda^{(\sigma)}\coloneqq\{\mu^{(\sigma)}\}$ specified by permutations $\sigma$ of order $n$, given by permutation densities $p_{\sigma}$ pushed forward coordinate-wise using the (left-continuous) inverse CDF $F_{\eta}^{-1}$. Suppose also that the marginal measure $\eta$ is absolutely continuous with respect to Lebesgue measure. Consider also a continuous function $f:\mathbb{R}\to\mathbb{R}$ whose expectation is to be estimated using $m=2$ coupled samples. In the limit $n\to\infty$, the class $\Lambda^{(\sigma)}$ contains measures that converge weakly to the optimal transport plan – that is, the sample coupling which provides the smallest possible estimator variance.
@@
$\mu^{*}=\textrm{arg min}_{\mu\in\Lambda_{2}(\eta)}\left[\mathbb{E}_{(\omega_{1},\omega_{2})\sim\mu}f(\omega_{1})f(\omega_{2})\right].$ (69)
@@
$\mu_{n}^{*}=\textrm{arg min}_{\mu\in\Lambda_{2}(\eta_{n})}\left[\mathbb{E}_{(\omega_{1},\omega_{2})\sim\mu}f(\omega_{1})f(\omega_{2})\right].$ (70)
@@
The OT plan is (asymptotically) in our search class. Lastly, we have that our search class of couplings $\Lambda^{(\sigma)}$ (corresponding to permutation densities pushed forward by $F_{\eta}^{-1}$) contains measures that converge in distribution to $\mu_{n}^{(\sigma)}$ when $n\to\infty$. In this mathematical sense, the asymptotic limit the class of couplings amongst which we optimise includes measures that give the greatest possible variance reduction: roughly speaking, our method is ‘asymptotically optimal’. This is intuitive because as the order of the permutation grows each ‘tile’ of nonzero density narrows and can be increasingly well-approximated by a delta function.
@@ E.1 Setup: estimating PageRank with random walks
$\widetilde{\mathbf{P}}\coloneqq(1-p_{\textrm{halt}})\mathbf{P}+\frac{p_{\textrm{halt}}}{N}\mathbf{E}.$ (75)
Here, $p_{\textrm{halt}}\in(0,1)$ is a scalar, $N$ is the number of nodes and $\mathbf{E}=[1]_{i,j\in\mathcal{N}}$ is a matrix whose entries are all ones. $\mathbf{P}$ is the transiton matrix of a simple random walk,
@@
where we normalised the sum of vector entries to $1$. Computing $\bm{\rho}$ analytically is expensive. Rearranging and Taylor expanding $(1-(1-p_{\textrm{halt}})\mathbf{P})^{-1}$, it is straightforward to see that
$\bm{\rho}_{i}=\frac{p_{\textrm{halt}}}{N}\sum_{j\in\mathcal{N}}\sum_{k=0}^{\infty}(1-p_{\textrm{halt}})^{k}\mathbf{P}^{k}_{ji}.$ (78)
This is simply a sum over all walks from each of the graph nodes $v_{j}$ to node $v_{i}$, weighted by their respective probabilities – that is, the expected number of random walkers ending at node $v_{i}$ if they terminate with probability $p_{\textrm{halt}}$ at every timestep – which invites the estimator proposed by Fogaras et al. [2005],
@@ E.2 OT formulation of PageRank variance reduction
$\end{multlined}\mathbb{E}(\widehat{\bm{\rho}}_{i}^{2})=\mathbb{E}\bigg{[}\frac{1}{N^{2}m^{2}}\sum_{v_{j},v_{k}\in\mathcal{N}}\sum_{n_{1},n_{2}=1}^{2}\mathbb{I}[n_{1}\textrm{th walk from node }v_{j}\textrm{ terminates at node}\hskip 2.84526ptv_{i}]\bigg{.}\\ \bigg{.}\cdot\mathbb{I}[n_{2}\textrm{th walk from node }v_{k}\textrm{ terminates at node}\hskip 2.84526ptv_{i}]\bigg{]}.$ (80)
@@ E.3 Empirical results for PageRank
Figure 9: PageRank estimator errors $\|\bm{\rho}-\widehat{\bm{\rho}}\|_{2}$ for a range of real-world graphs when $2$ random walkers are taken to be i) i.i.d., ii) antithetic or iii) $\sigma$-coupled. Lower is better. The error is always at least as small and often substantially better when the walkers are coupled using our OT coupling. One standard error is shaded.
We now provide experiments using our OT coupling to improve the convergence of PageRank estimates. Fig. 9 shows the results. For termination probabilities $p_{\textrm{halt}}\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$, we compute the optimal permutation $\sigma$ by solving the matching problem in Eq. 82 on an Erdős-Rènyi graph with $N=100$ nodes, taking permutation order $n=10$ as a hyperparameter. We then test these couplings on a range of real-world graphs [Ivashkin, 2023] of different sizes, plotting the estimator error $\|\bm{\rho}-\widehat{\bm{\rho}}\|_{2}$ (where $\bm{\rho}$ is the true PageRank vector and $\widehat{\bm{\rho}}$ is its MC estimate) against the termination probability $p_{\textrm{halt}}$. We include walkers that are i.i.d., antithetic [Reid et al., 2024c] and $\sigma$-coupled. At every value of $p_{\textrm{halt}}$, our method performs at least as well as the baselines, and in some cases (e.g. on the very large cora graph) it gives much greater variance reduction.
Note that we only consider connected subgraphs and take $1000$ repeats for standard errors. Reading the plots from left to right for the top row and then the bottom, the number of nodes are: $\{34,62,105,115,398,986,1133,1272,2120,2708\}$. The shape of the curve and size of the gain from coupling depends on the particular graph being considered. For some graphs it is not possible to obtain a big reduction in PageRank estimator variance by coupling walkers (e.g. football and newsgroup), so neither antithetic termination nor our $\sigma$-coupling can provide a big gain. A detailed investigation of how these observations relate to the mathematical graph structure is deferred to future work.
@@ F.1 More results for Gram matrix estimation
First, we run the experiment in the first part of Sec. 4.2 for more real-world graphs. We approximate the $2$-regularised Laplacian kernel $\smash{(\mathbf{I}-\sigma^{2}\widetilde{\mathbf{L}})^{-2}}$ with a regularisation parameter $\sigma=1$ and $m=2$ walkers. Fig. 10 shows the results, with the kernel approximation error normalised by the i.i.d. result (unlike in Fig. 3) for clarity of comparison. The $\sigma$-coupling consistently reduces the estimator variance, providing equally good or better results compared to the hand-crafted antithetic termination algorithm. The size of the gain depends on the graph structure.
Figure 10: Relative Frobenius norm error $\|\mathbf{K}-\widehat{\mathbf{K}}\|_{\textrm{F}}/\|\mathbf{K}\|_{\textrm{F}}$ between true and approximated Gram matrices for different termination probabilities when walkers are i) i.i.d., ii) antithetic [Reid et al., 2024c] or iii) $\sigma$-coupled using our OT approach. All results are normalised by the i.i.d. variance for easy comparison. Lower is better. The $\sigma$-coupling is consistently as good or better across values of $p_{\textrm{halt}}$ and graph topologies and sizes. One standard error is shaded.
@@ F.2 Scalable GPs on graphs
Like their Euclidean counterparts, exact GPs on graphs suffer $\mathcal{O}(N^{3})$ time complexity, making them impractical for very large structures. Techniques to improve their scalability include graph Fourier feature approximations, which approximate the kernel matrix using a truncated eigenvalue expansion computed using the Lanczos algorithm.  The price of increased speed is that our kernel estimate is biased. Another approach is to restrict to kernels whose inverses are sparse, e.g. a Matérn kernel with a small integer $\nu$-parameter, at the expense of lost flexibility [Borovitskiy et al., 2021]. In this paper, we have proposed to instead use GRFs, which give an unbiased estimate to the kernel matrix in subquadratic time. GRFs are sparse: intuitively, $\phi_{\textrm{GRF}}(v_{i})$ is constructed by simulating $m$ random walks out of $v_{i}$ and adding weighted contributions only at the coordinates corresponding to visited nodes. For walks with geometrically-distributed lengths this is typically a small subset of $\mathcal{N}$. Therefore, the matrix $\smash{\widehat{\mathbf{K}}\coloneqq[\phi(v_{i})^{\top}\phi(v_{j})]_{i,j=1}^{N}}$ is a sparse, unbiased estimate to any kernel, not just specific families. [^4] Strictly: any graph node kernel that is expressible as a function of a weighted adjacency matrix. We can use established numerical routines for sparse linear algebra to speed up training, providing computational savings on big graphs. This is analogous to the sparse spectrum GPs discussed in App. B [Lázaro-Gredilla et al., 2010]; the interested reader might benefit from reviewing this section first.
@@
Training GRF-GPs: technical details. Here, provide details to supplement Sec. 4.2. We use mesh graphs made available by Dawson-Haggerty [2023], taking the faces as our graph nodes. Reading left to right, the number of nodes are: $894,999,1280,1572,3838,8700$. For probabilistic mesh interpolation, we use the diffusion kernel $\mathbf{K}=\kappa\exp(-\gamma^{2}\widetilde{\mathbf{L}})$. Here, $\widetilde{\mathbf{L}}$ is the normalised graph Laplacian (Eq. 48) and $\kappa,\gamma$ are learnable parameters corresponding to the signal variance and kernel lengthscale respectively. The observation noise $\sigma$ is also learnable. For every graph, we construct a kernel estimate by sampling $m=8$ i.i.d. random walks out of every node. We then train the corresponding vanilla GP by optimising the log marginal likelihood on observed nodes, using the Adam optimiser [Kingma and Ba, 2014] for $1000$ epochs (after which both the objective value and the hyperparameters are empirically found to have converged). We freeze $\kappa,\gamma,\sigma$ and compute the corresponding exact kernel. We approximate this exact kernel using $\{16,32,64\}$ random walks that are either i.i.d., antithetic or $\sigma$-coupled. We perform approximate Bayesian inference in each case, computing the accuracy of the kernel estimate, the average test RMSE and the KL divergence to the true posterior distribution (unnormalised by the number of test points). We use permutations of order $n=50$ and a termination probability $p_{\textrm{halt}}=0.4$. As reported in the main body, coupling the lengths of graph random walkers permits better approximate Bayesian inference with GPs on graphs.
@@ F.3 Probabilistic traffic interpolation
Figure 11: Results for probabilistic traffic interpolation experiment. Plots show the kernel approximation accuracy, test RMSE and KL divergence to the true posterior for graph kernels estimated using i.i.d., antithetic and $\sigma$-coupled GRFs. Lower is better. $\sigma$-coupling gives the best results for approximate inference. One standard error over random draws of walkers is shaded.